OpenStudy (anonymous):

solve. x= log[3] 1/243

OpenStudy (anonymous):

x = -5

OpenStudy (anonymous):

1/243 = 3 ^ x 1/243 = 3^ -5 x = -5

OpenStudy (alfie):

3^x = 1/243 3^x = 1/3^5 => x= -5